# eKalibr-Stereo: Continuous-Time Spatiotemporal Calibration for Event-Based Stereo Visual Systems

### v1

@@ Abstract
The bioinspired event camera, distinguished by its exceptional temporal resolution, high dynamic range, and low power consumption, has been extensively studied in recent years for motion estimation, robotic perception, and object detection. In ego-motion estimation, the stereo event camera setup is commonly adopted due to its direct scale perception and depth recovery. For optimal stereo visual fusion, accurate spatiotemporal (extrinsic and temporal) calibration is required. Considering that few stereo visual calibrators orienting to event cameras exist, based on our previous work eKalibr (an event camera intrinsic calibrator), we propose eKalibr-Stereo for accurate spatiotemporal calibration of event-based stereo visual systems. To improve the continuity of grid pattern tracking, building upon the grid pattern recognition method in eKalibr, an additional motion prior-based tracking module is designed in eKalibr-Stereo to track incomplete grid patterns. Based on tracked grid patterns, a two-step initialization procedure is performed to recover initial guesses of piece-wise B-splines and spatiotemporal parameters, followed by a continuous-time batch bundle adjustment to refine the initialized states to optimal ones. The results of extensive real-world experiments show that eKalibr-Stereo can achieve accurate event-based stereo spatiotemporal calibration. The implementation of eKalibr-Stereo is open-sourced at (https://github.com/Unsigned-Long/eKalibr) to benefit the research community.
@@ I Introduction and Related Works
Stereo visual spatiotemporal calibration typically consists of two sub-modules: ($i$) correspondence construction (front end) and ($ii$) spatiotemporal optimization (back end). In the front end, artificial visual targets, such as checkerboards [5], April Tags [6], and ChArUco board [7], are commonly employed to construct accurate 3D-2D correspondences with real-world geometric scale through pattern recognition. While a substantial number of target pattern recognition methods [5, 8, 9] oriented to standard cameras have been proposed, they are not applicable to event cameras, which output asynchronous event stream rather than conventional intensity images. To recognize target patterns from raw events, early works [10, 11, 12] generally rely on blinking light emitting diode (LED) grid boards. Although target patterns can be accurately extracted, requiring additional LED boards introduces inconvenience. Meanwhile, these methods typically require the event camera to remain stationary, making them unsuitable for multi-camera spatiotemporal calibration that necessitates motion excitation [13]. To address this, subsequent methods [14, 15] have proposed an alternative approach, namely reconstructing intensity images from raw events using event-based image reconstruction methods (such as E2VID[16] and Spade-E2VID [17]) first, followed by conventional image-based pattern recognition methods. Although reconstructed images exhibit high consistency, substantial noise within the images could lead to imprecise pattern extraction, which further affects calibration accuracy. Considering these, some event-based pattern recognition methods have been proposed recently, aiming to extract target patterns from dynamically acquired raw events directly. Based on density-based spatial clustering (DBSCAN), Huang et al. [18] cluster events accumulated in a time window and fit circles to extract centers of circles on a circle grid board. Similarly but performing DBSCAN in spatiotemporal domain, Salah et al. [19] cluster events generated from a circle grid board, and then fit 3D cylinders for circle center determination. Focusing on event-frame stereo camera calibration, Wang et al. [20] designed a novel circle grid board with cross points, and match candidate event regions of opposite polarity for 3D oblique elliptical cylinder fitting. In our previous work eKalibr [21], an event-only target pattern recognition method oriented to the circle grid board is proposed for intrinsic calibration, which cluster events and match clustering pair based on normal flow estimation, and fit 3D elliptical cylinders for circle center extraction.
@@
In this article, focusing on event-based stereo visual systems, we present a continuous-time-based spatiotemporal calibration method, named eKalibr-Stereo, to accurately estimate the extrinsics and time offset between event cameras. Building upon our previous work eKalibr [21], eKalibr-Stereo tracks continuous circle grid patterns (complete and incomplete ones) from raw events for 3D-2D correspondence construction. Given the high non-linearity of continuous-time optimization, a two-stage initialization procedure is first conducted to recover the initials of states, which are then iteratively refined to optimal ones using a continuous-time-based batch bundle adjustment. eKalibr-Stereo makes the following (potential) contributions:
@@
    We proposed a continuous-time-based spatial and temporal calibrator for event-based stereo visual systems, which could accurately determine both extrinsics and time offset of a stereo event camera system.
@@
Note that the proposed eKalibr-Stereo supports one-shot event-based multi-camera spatiotemporal calibration (an arbitrary number of event cameras). To enhance clarity, this article only considers the minimal stereo event camera configuration, as it’s the most critical sensor setup for facilitating multi-camera calibration.
@@ II-A Notations and Definitions
Given a raw event $\boldsymbol{\mathrm{e}}$ generated by the event camera, we use $\tau\in\mathbb{R}$, $\boldsymbol{\mathrm{x}}\in\mathbb{Z}^{2}$, and $p\in\{{\text{-}}1,{\text{+}}1\}$ to represent its timestamp, pixel position, and polarity respectively, i.e., $\boldsymbol{\mathrm{e}}\triangleq\{\tau,\boldsymbol{\mathrm{x}},p\}$. The camera frame and world frame (defined by the circle grid board) are represented as $\underrightarrow{\mathcal{F}}_{c}$ and $\underrightarrow{\mathcal{F}}_{w}$, respectively. The transformation from $\underrightarrow{\mathcal{F}}_{c}$ to $\underrightarrow{\mathcal{F}}_{w}$ are parameterized as the Euclidean matrix ${\boldsymbol{\mathrm{T}}_{c}^{w}}\in\mathrm{SE(3)}$, which is defined as:
@@ III Methodology
This section details the proposed event-based continuous-time stereo visual spatiotemporal calibrator eKalibr-Stereo.
@@ III-A System Overview
The comprehensive framework of the proposed event-based stereo visual calibrator is illustrated in Fig. 2. Given raw asynchronous event streams from the stereo camera rig, we first perform normal flow estimation and ellipse fitting for each event camera, to track both complete and incomplete circle grid patterns. The event camera that tracks the grid pattern the most among two cameras would be treated as the reference (primary) camera (denoted as $\underrightarrow{\mathcal{F}}_{c_{r}}$), while the other would be assigned as the target (secondary) camera (denoted as $\underrightarrow{\mathcal{F}}_{c_{t}}$). Subsequently, for each tracked grid pattern of the reference camera, PnP [35] would be employed to estimate the camera pose based on 3D-2D correspondences. Discrete poses of the reference camera are then utilized to recover a continuous-time trajectory. The spatiotemporal parameters, i.e., extrinsics and time offset, would be also initialized based on continuous-time hand-eye alignment. Finally, a continuous-time-based visual bundle adjustment would be performed to refine all states in the estimator to the global optimal ones.
@@ III-B Event-Based Circle Grid Tracking
Given generated raw event streams, we first employ the event-based circle grid pattern recognition algorithm [21] proposed in our previous work eKalibr to extract complete grid patterns for each camera. As described in [21], we first perform event-based normal flow estimation [36] on the surface of active event (SAE) [37] and homopolarly cluster inlier events for cluster matching. Spatiotemporal ellipses would then be estimated for each matched cluster pair for center determination of the grid circle. Finally, temporally synchronized centers would be organized as ordered grid patterns. Note that although both the asymmetric and symmetric circle grids are supported in eKalibr, the asymmetric circle grid is utilized in this work as the visual target, as it does not exhibit 180-degree ambiguity [38].
Figure 3: Schematic of incomplete grid pattern tracking. When ellipses corresponding to all gird circles are fitted (subfigure A), a complete pattern (subfigure B) can be extracted. However, due to oblique perspective (subfigure C), noisy events (subfigure D), or the grid being outside the sensing range (subfigure E), not all ellipses can be identified, resulting in incomplete pattern.
@@ III-C2 Spatiotemporal Initialization
After the continuous-time trajectories of the reference camera are initialized, we employ the continuous-time hand-eye alignment to initialize the extrinsics and time offset of the target camera with respect to the reference camera. This could be achieved by solving the following least-squares problem:
@@ IV Real-World Experiment
IV-A Experiment Setup
Fig. 4 shows the self-assembled sensor rig for real-world experiments, consisting of two software-synchronized DAVIS346 event cameras (the resolution is 346$\times$260). We refer to the two event cameras as the left camera and the right camera for convenience in subsequent description and discussion. To ensure the comprehensiveness of the experiment, three different sizes of asymmetric circle grid patterns (3$\times$7, 4$\times$9, and 4$\times$11), as shown in Fig. 4, are used in real-world experiments. The radius rate and spacing for all grid boards are 2.5 and 50 mm, respectively. Altogether 5 sequences of 30-second data are collected for each grid board for Monte-Carlo evaluation.
To ensure the reliability of temporal calibration evaluation of eKalibr-Stereo in experiments, we manually shifted the timestamps of all events generated by the right camera by 50 ms after data acquisition, to simulate a stereo visual system with a time offset, i.e., we have $\tau_{\mathrm{right}}\leftarrow\tau_{\mathrm{right}}-\Delta\tau_{\mathrm{shift}}$ where $\Delta\tau_{\mathrm{shift}}=0.05$ sec. Therefore, the time offset of the right camera with respect to the left camera, i.e., ${\tau_{\mathrm{right}}^{\mathrm{left}}}$, is theoretically equal to 0.05 sec.
@@ IV-B Evaluation and Comparison of Grid Pattern Tracking
Figure 5: Grid tracking performance of eKalibr-Stereo for three different sizes of ACircle grid boards. Based on the time distance threshold of neighbor patterns (set $\Delta\tau_{\mathrm{thd}}$ as 0.1 sec here) and the continuity threshold of grid tracking (set $\mathcal{N}_{\mathrm{thd}}$ as 50 here), tracked grids are segmented and corresponding piece-wise B-splines would be constructed. Details can be found in Section III-C1.
TABLE I: Evaluation and Comparison of Circle Gird Tracking
@@
    The tracking success rate is obtained by: the number of successful grid trackings divided by the total number of grids to track.
@@ IV-C Evaluation and Comparison of Calibration Performance
To comprehensively and quantitatively evaluate the spatiotemporal calibration performance of the proposed eKalibr-Stereo, we conducted real-world Monte-Carlo experiments. A total of three stereo visual calibration methods were incorporated in experiments for the evaluation and comparison of calibration results:
@@
    Frame-Based Stereo Calibration (DV Software [42]): The stereo frame-based (standard) visual extrinsic calibration toolkit provided by iniVation, i.e., the developer of DAVIS346 event camera employed in our real-world experiments. Since the DAVIS346 event camera supports standard frame output, stereo visual extrinsics can be accurately determined using the conventional frame-oriented calibration pipeline in DV Software. Therefore, the calibration results from DV Software can be treated as the reference. Note that DV Software only supports spatial (extrinsic) calibration, requiring the stereo event camera rig to be temporally synchronized.
@@
    Event-Only eKalibr-Stereo: The proposed event-based stereo visual calibrator, supports both spatial (extrinsics) and temporal (time offset) determination.
@@
Method Grid Type Extrinsic Temporal
@@
Roll Pitch Yaw X Y Z Est. Ref.
DV Software [42]    3×$\times 7$0.47$\pm$0.01 0.70$\pm$0.02 0.38$\pm$0.02 12.07$\pm$0.01 -0.01$\pm$0.01 -0.38$\pm$0.04 50.00
   4×$\times 9$0.42$\pm$0.02 0.69$\pm$0.01 0.38$\pm$0.02 12.08$\pm$0.02 -0.01$\pm$0.01 -0.40$\pm$0.05 50.00
   4×$\times 11$0.45$\pm$0.01 0.68$\pm$0.02 0.39$\pm$0.01 12.07$\pm$0.02 -0.01$\pm$0.01 -0.39$\pm$0.05 50.00
E2VID [16] & Kalibr [25]    3×$\times 7$1.03$\pm$1.12 1.96$\pm$1.34 1.26$\pm$1.00 13.75$\pm$1.58 2.01$\pm$1.44 -2.23$\pm$1.39 57.37$\pm$4.48 50.00
   4×$\times 9$1.37$\pm$1.42 1.74$\pm$1.60 1.06$\pm$1.28 11.23$\pm$1.68 3.42$\pm$1.51 -1.72$\pm$1.25 55.92$\pm$4.03 50.00
   4×$\times 11$0.95$\pm$1.16 1.52$\pm$1.33 0.81$\pm$1.01 11.77$\pm$1.46 1.82$\pm$1.23 -1.55$\pm$1.19 54.57$\pm$3.99 50.00
eKalibr-Stereo    3×$\times 7$0.48$\pm$0.03 0.71$\pm$0.02 0.39$\pm$0.01 12.08$\pm$0.01 -0.01$\pm$0.01 -0.36$\pm$0.06 49.58$\pm$0.10 50.00
   4×$\times 9$0.36$\pm$0.01 0.67$\pm$0.02 0.38$\pm$0.01 12.07$\pm$0.01 -0.01$\pm$0.02 -0.42$\pm$0.06 49.98$\pm$0.06 50.00
   4×$\times 11$0.43$\pm$0.01 0.66$\pm$0.01 0.39$\pm$0.01 12.07$\pm$0.01 -0.01$\pm$0.01 -0.37$\pm$0.08 49.93$\pm$0.06 50.00
@@
    The value in each table cell is represented as (Estimate Mean) $\pm$ (STD).
Table II summarized final spatiotemporal calibration results of DV Software, E2VID & Kalibr, and eKalibr-Stereo in real-world Monte-Carlo experiments, showing the spatiotemporal estimates and corresponding standard deviations (STDs). As can be seen, the calibration method using E2VID and Kalibr achieved the poorest results with the largest STDs and biases compared with the frame-based DV Software. This is mainly due to the high noise in reconstructed image frames from E2VID (though reconstructed images are consistent on a macroscopic scale), which could reduce the extraction accuracy of the grid board in kalibr, further affecting the spatiotemporal determination. In comparison, eKalibr-Stereo is able to directly and accurately extract grid patterns from raw events, thereby achieving calibration results comparable to frame-based DV Software using continuous-time spatiotemporal optimization. Overall, eKalibr-Stereo can achieve an average STD of less than 0.05 degrees for extrinsic rotation determination, 0.05 cm for extrinsic translation, and 0.1 ms for time offset.
Figure 6: The distributions of projection errors after spatiotemporal optimization for two cameras in one solving. The subplots on the left correspond to the left camera, while the subplots on the right correspond to the right camera.
Fig. 6 plotted the distributions of projection errors of two cameras after spatiotemporal calibration using three different sizes of grid boards as visual targets in eKalibr-Stereo. It can be found that the projection errors follow a zero-mean normal distribution, indicating that the calibrated spatiotemporal parameters are well-estimated and unbiased. The average sigma of final projection errors is less than 0.1 pixels, indicating the high-accuracy spatiotemporal calibration eKalibr-Stereo capable of.

### v2

@@ Abstract
The bioinspired event camera, distinguished by its exceptional temporal resolution, high dynamic range, and low power consumption, has been extensively studied in recent years for motion estimation, robotic perception, and object detection. In ego-motion estimation, the stereo event camera setup is commonly adopted due to its direct scale perception and depth recovery. For optimal stereo visual fusion, accurate spatiotemporal (extrinsic and temporal) calibration is required. In this letter, we present eKalibr-Stereo, an accurate spatiotemporal calibrator for event-based stereo visual systems, utilizing the widely used circular grid board. To ensure the continuity of grid pattern tracking, building upon the grid pattern recognition method in eKalibr, an additional motion prior-based tracking module is designed in eKalibr-Stereo to track incomplete grid patterns. Based on tracked grid patterns, a two-step initialization procedure is performed to recover initial guesses of piece-wise B-splines and spatiotemporal parameters, followed by a continuous-time batch bundle adjustment to refine the initialized states to optimal ones. The results of extensive real-world experiments show that eKalibr-Stereo can achieve accurate event-based stereo spatiotemporal calibration. The implementation of eKalibr-Stereo is open-sourced at (https://github.com/Unsigned-Long/eKalibr) to benefit the research community.
@@ I Introduction and Related Works
Stereo visual spatiotemporal calibration typically consists of two sub-modules: ($i$) correspondence construction (front end) and ($ii$) spatiotemporal optimization (back end). In the front end, artificial visual targets, such as checkerboards [5], April Tags [6], and ChArUco board [7], are commonly employed to construct accurate 3D-2D correspondences with real-world geometric scale through pattern recognition. While a substantial number of target pattern recognition methods [5, 8, 9] oriented to standard cameras have been proposed, they are not applicable to event cameras, which output asynchronous event stream rather than conventional intensity images. To recognize target patterns from raw events, early works [10, 11, 12] generally rely on blinking light emitting diode (LED) grid boards. Although target patterns can be accurately extracted, requiring additional LED boards introduces inconvenience. Meanwhile, these methods typically require the event camera to remain stationary, making them unsuitable for multi-camera spatiotemporal calibration that necessitates motion excitation [13]. To address this, subsequent methods [14, 15] have proposed an alternative approach, namely reconstructing intensity images from raw events using event-based image reconstruction methods (such as E2VID[16] and Spade-E2VID [17]) first, followed by conventional image-based pattern recognition methods. Although reconstructed images exhibit high consistency, substantial noise within the images could lead to imprecise pattern extraction, which further affects calibration accuracy. Considering these, some event-based pattern recognition methods have been proposed recently, aiming to extract target patterns from dynamically acquired raw events directly. Based on density-based spatial clustering (DBSCAN), Huang et al. [18] cluster events accumulated in a time window and fit circles to extract centers of circles on a circle grid board. Similarly but performing DBSCAN in spatiotemporal domain, Salah et al. [19] cluster events generated from a circle grid board, and then fit 3D cylinders for circle center determination. DBSCAN-based clustering methods are sensitive to hyperparameters, which can lead to instability in grid pattern recognition and further affect the calibration. Differently and more directly, the authors of EF-Calib [20] specialized in target design and developed a novel circular grid board with cross points, which enables efficient and accurate event-based (using circle edges) and frame-based (using cross points) target recognition for subsequent event-frame stereo camera calibration. In this method, events of circle edges are clustered using BBDT and then matched based on cluster distance. Similarly, the authors of the event-based intrinsic calibrator eKalibr [21] proposed an event-only target pattern recognition method designed for the commonly used circular grid board. This method clusters and matches event clusters based on normal flow estimation (from first principles of events), rather than relying on traditional algorithms like BBDT in [20], resulting in more efficient and explainable extraction.
@@
In this article, focusing on event-based stereo visual systems, we present a continuous-time-based spatiotemporal calibration method, named eKalibr-Stereo, to accurately estimate the extrinsics and time offset between event cameras. Building upon eKalibr [21], eKalibr-Stereo tracks continuous circle grid patterns (complete and incomplete ones) from raw events for 3D-2D correspondence construction. Given the high non-linearity of continuous-time optimization, a two-stage initialization procedure is first conducted to recover the initials of states, which are then iteratively refined to optimal ones using a continuous-time-based batch bundle adjustment. eKalibr-Stereo makes the following (potential) contributions:
@@
    We proposed a continuous-time-based spatial and temporal calibrator for event-based stereo visual systems, which could accurately determine both extrinsics and time offset of a stereo event camera system. To the best of our knowledge, this is the first open-source work focused on event-based stereo spatiotemporal calibration.
@@
Note that the proposed eKalibr-Stereo supports one-shot event-based multi-camera spatiotemporal calibration (an arbitrary number of event cameras). To enhance clarity, this article only considers the minimal stereo event camera configuration, as it’s the most typical sensor setup for facilitating multi-camera calibration.
@@ II-A Notations and Definitions
Given a raw event $\boldsymbol{\mathrm{e}}$ generated by the event camera, we use $\tau\in\mathbb{R}$, $\boldsymbol{\mathrm{x}}\in\mathbb{Z}^{2}$, and $p\in\{{\text{-}}1,{\text{+}}1\}$ to represent its timestamp, pixel position, and polarity, respectively, i.e., $\boldsymbol{\mathrm{e}}\triangleq\{\tau,\boldsymbol{\mathrm{x}},p\}$. The camera frame and world frame (defined by the circle grid board) are represented as $\underrightarrow{\mathcal{F}}_{c}$ and $\underrightarrow{\mathcal{F}}_{w}$, respectively. The transformation from $\underrightarrow{\mathcal{F}}_{c}$ to $\underrightarrow{\mathcal{F}}_{w}$ are parameterized as the Euclidean matrix ${\boldsymbol{\mathrm{T}}_{c}^{w}}\in\mathrm{SE(3)}$, which is defined as:
@@ III Methodology
This section presents the proposed event-based continuous-time stereo visual spatiotemporal calibration framework.
@@ III-A System Overview
The comprehensive framework of the proposed event-based stereo visual calibrator is illustrated in Fig. 2. Given raw asynchronous event streams from the stereo camera rig, we first perform normal flow estimation and ellipse fitting for each event camera, to track both complete and incomplete circle grid patterns, see Section III-B. The event camera that tracks the grid pattern the most among two cameras would be treated as the reference (primary) camera (denoted as $\underrightarrow{\mathcal{F}}_{c_{r}}$), while the other would be assigned as the target (secondary) camera (denoted as $\underrightarrow{\mathcal{F}}_{c_{t}}$). Subsequently, for each tracked grid pattern of the reference camera, PnP [35] would be employed to estimate the camera pose based on 3D-2D correspondences. Discrete poses of the reference camera are then utilized to recover a continuous-time trajectory, see Section III-C1. The spatiotemporal parameters, i.e., extrinsics and time offset, would be also initialized based on continuous-time hand-eye alignment, see Section III-C2. Finally, a continuous-time-based visual bundle adjustment would be performed to refine all states in the estimator to the global optimal ones, see Section III-D.
@@ III-B Event-Based Circle Grid Tracking
Given generated raw event streams, we first employ the event-based circle grid pattern recognition algorithm [21] proposed in eKalibr to extract complete grid patterns for each camera. As described in [21], we first perform event-based normal flow estimation [36] on the surface of active event (SAE) [37] and homopolarly cluster inlier events for cluster matching. Spatiotemporal ellipses would then be estimated for each matched cluster pair for center determination of the grid circle. Finally, temporally synchronized centers would be organized as ordered grid patterns. Note that although both the asymmetric and symmetric circle grids are supported in eKalibr, the asymmetric circle grid is utilized in this work, as it does not exhibit 180-degree ambiguity [38].
Figure 3: Schematic of incomplete grid pattern tracking. When ellipses corresponding to all grid circles are fitted (subfigure A), a complete pattern (subfigure B) can be extracted. However, due to oblique perspective (subfigure C), noisy events (subfigure D), or the grid being outside the sensing range (subfigure E), not all ellipses can be identified, resulting in incomplete pattern.
@@ III-C2 Spatiotemporal Initialization
After the continuous-time trajectories of the reference camera are initialized, we employ the continuous-time hand-eye alignment [^2] Hand-eye alignment refers to the process of recovering the spatiotemporal parameters between different sensors through aligning their kinematics based on rigid-body constraints. [13] to initialize the extrinsics and time offset of the target camera with respect to the reference camera. This could be achieved by solving the following least-squares problem:
@@ IV Real-World Experiment
IV-A Equipment Setup
Fig. 4 shows the self-assembled sensor suite for real-world experiments, consisting of two hardware-synchronized DAVIS346 event cameras (the resolution is 346$\times$260). We refer to the two event cameras as the left camera and the right camera for convenience in subsequent description and discussion. To ensure the comprehensiveness of the experiment, three different sizes of asymmetric circle grid patterns (3$\times$7, 4$\times$9, and 4$\times$11), as shown in Fig. 4, are used in real-world experiments. The radius rate and spacing for all grid boards are 2.5 and 50 mm, respectively.
@@ IV-B Evaluation and Comparison of Grid Pattern Tracking
Figure 5: Grid tracking performance of eKalibr-Stereo for three different sizes of ACircle grid boards. Based on the time distance threshold $\Delta\tau_{\mathrm{thd}}$ (set to 0.1 sec here) and the continuity threshold $\mathcal{N}_{\mathrm{thd}}$ (set to 50 here), tracked grids are segmented (black lines) and corresponding piece-wise B-splines would be constructed. Note that the total span of the segments is related to the tracking continuity, rather than the board size. More details are shown in Section III-C1.
TABLE I: Evaluation and Comparison of Circle Grid Tracking
@@
    Cmp. Grid and Incmp. Grid refer to complete grids and incomplete grids, respectively. The tracking success rate is the ratio of successful grid trackings to the total grids.
@@ IV-C Evaluation and Comparison of Calibration Performance
To comprehensively and quantitatively evaluate the spatiotemporal calibration performance of the proposed eKalibr-Stereo, we conducted real-world Monte-Carlo experiments, where 5 sequences of 30-second data are collected for each grid board for evaluation. A total of four stereo visual calibration methods were incorporated in experiments for the evaluation and comparison of calibration results:
@@
    Frame-Based Stereo Calibration (DV Software [42]): The stereo frame-based (standard) visual extrinsic calibration toolkit provided by iniVation, i.e., the developer of DAVIS346 event camera employed in our real-world experiments. Since the DAVIS346 event camera supports standard frame output, stereo visual extrinsics can be accurately determined using the conventional frame-oriented calibration pipeline in DV Software. Therefore, the calibration results can be treated as the reference.
@@
    eKalibr-Stereo without Incomplete Grid Tracking: The proposed event-based stereo visual calibrator, supports both spatial (extrinsics) and temporal (time offset) determination. Note that only the extracted complete grids are used for solving in this method.
  4. 4.
    eKalibr-Stereo with Incomplete Grid Tracking: The proposed event-based stereo spatiotemporal calibrator. Note that both the extracted complete grids and tracked incomplete grids are used for solving in this method.
Meanwhile, to ensure the evaluation of temporal calibration of eKalibr-Stereo in experiments, we manually shifted the timestamps of all events generated by the right camera by $\Delta\tau_{\mathrm{shift}}$ after data acquisition, to simulate a stereo visual system with a time offset, i.e., we have $\tau_{\mathrm{right}}\leftarrow\tau_{\mathrm{right}}-\Delta\tau_{\mathrm{shift}}$. Therefore, the time offset of the right camera with respect to the left camera, i.e., ${\tau_{\mathrm{right}}^{\mathrm{left}}}$, is theoretically equal to $\Delta\tau_{\mathrm{shift}}$. $\Delta\tau_{\mathrm{shift}}$ is sequentially set to various values in the experiment, namely: 10 ms, 20 ms, 50 ms, and 100 ms, for a comprehensive evaluation. Note that DV Software only supports spatial (extrinsic) calibration, requiring the stereo event camera rig to be temporally synchronized, and thus are not considered in time offset evaluation.
@@
Extrinsic Temporal
@@
Method Roll Pitch Yaw X Y Z Est. Ref.
[…]
@@
    The value in each table cell is represented as (Estimate Mean) $\pm$ (STD). A smaller STD indicates better repeatability and stability of the method.
    The best results are highlighted in bold, while the second-best results are underlined, under the condition where the extrinsics from DV Software and the simulated time delays are used as references (i.e., values with gray background).
Table II summarized final spatiotemporal calibration results of DV Software, E2VID & Kalibr, and eKalibr-Stereo (w/o and w/ Incmp. Grid) in real-world Monte-Carlo experiments, showing the spatiotemporal estimates and corresponding standard deviations (STDs). As can be seen, the calibration method using E2VID & Kalibr achieved the poorest results with the largest STDs and biases. Compared to the values from frame-based DV Software, the estimated values from E2VID & Kalibr deviated by approximately 0.3 degrees for extrinsic rotation (with a STD of 0.3 degrees), 1.0 cm for extrinsic translation (with a STD of 1.2 cm), and 3.0 ms for time offset (with a STD of 2.3 ms). This is mainly due to the high noise in reconstructed image frames from E2VID (though reconstructed images are consistent on a macroscopic scale), which could reduce the extraction accuracy of the grid board in Kalibr, further affecting the spatiotemporal determination. In comparison, eKalibr-Stereo is able to directly and accurately extract grid patterns from raw events, thereby achieving calibration results comparable to frame-based DV Software. Specifically, in the case of eKalibr-Stereo with and without incomplete grid tracking, eKalibr-Stereo with incomplete grid tracking is able to track more grid boards by utilizing motion priors, thereby providing stronger and more continuous motion constraints. As a result, it achieves better calibration results (closer to the estimates from DV Software) and improved repeatability (with a smaller STD), compared to eKalibr-Stereo without incomplete grid tracking. Overall, among the three event-based methods, eKalibr-Stereo with incomplete grid tracking achieves the best results, with the closest match to DV Software’s results and the smallest STD.
Figure 6: The distributions of projection errors after calibration. $\sigma$, $2\sigma$, and $3\sigma$ represent one, two, and three STDs of the reprojection errors, respectively.
Fig. 6 plotted the distributions of projection errors of two cameras after spatiotemporal calibration in eKalibr-Stereo. It can be found that the projection errors follow a zero-mean normal distribution, indicating that the calibrated spatiotemporal parameters are well-estimated and unbiased. The average sigma of final projection errors is less than 0.1 pixels, indicating the high-accuracy spatiotemporal calibration eKalibr-Stereo capable of.